10a. A string of length 1 m and linear density of 0.035 kg/m is stretched between 2 posts with a tension of 667 N. What is the frequency of the first 5 modes? Sketch these 5 modes. b. Now assume that you are plucking the string at a distance of 20 cm from one end. What will be the first four lowest harmonic frequencies? (Hint: You dont have to recalculate use the results from part a)
Answer #1

0% C Hering meet in give-y ว่า longth and m=1,2,3 S mce n- the -km da 2Lto = L | 667 2157.142 to= 69.023 ㎐ 73= 207 Hzt4 = 4×63= 276 Hz 4 = 274 Hz t5 = 5×63= 345 Hz- t5 = 345 Hz -The sketch 억 those 5 moda are mode ナ= 6gH2 Q = 138 H2 t3 = 207ana emal 160 5 Jongth (rom ens end) iumot have Pamonic siu be absent. e quem have been cal culaled h tha bart (a)